8-23 Variable manufacturing overhead variance analysis. The French Bread Company bakes baguettes for distribution to upscale grocery stores. The company has two direct-cost categories: direct materials and direct manufacturing labor. Variable manufacturing overhead is allocated to products on the basis of standard direct manufacturing labor-hours. Following is some budget data for the French Bread Company:

Direct manufacturing labor use

0.02 hours per baguette

Variable manufacturing overhead

$10.00 per direct manufacturing labor-hour

The French Bread Company provides the following additional data for the year ended December 31, 2017:

Planned (budgeted) output

3,200,000 baguettes

Actual production

2,800,000 baguettes

Direct manufacturing labor

50,400 hours

Actual variable manufacturing overhead

$680,400

Required:

1. What is the denominator level used for allocating variable manufacturing overhead? (That is, for how many direct manufacturing labor-hours is French Bread budgeting?)

2. Prepare a variance analysis of variable manufacturing overhead. Use Exhibit 8-4 (page 324) for reference.

3. Discuss the variances you have calculated and give possible explanations for them.

8-24 Fixed manufacturing overhead variance analysis (continuation of 8-23). The French Bread Company also allocates fixed manufacturing overhead to products on the basis of standard direct manufacturing labor-hours. For 2017, fixed manufacturing overhead was budgeted at $4.00 per direct manufacturing labor-hour. Actual fixed manufacturing overhead incurred during the year was $272,000.

Required:

1. Prepare a variance analysis of fixed manufacturing overhead cost. Use Exhibit 8-4 (page 324) as a guide.

2. Is fixed overhead underallocated or overallocated? By what amount?

3. Comment on your results. Discuss the variances and explain what may be driving them

plz do the 8-24 and use comupter writing
Answer #1

23- actual time = 50400

budgeted time = 560000

actual rate = 680400/50400 = 13.5

budgeted rate = 10

1) budgeting manufacturing labour hour = budgeting manufacturing labour use * actual production of baguettes

= 0.02 * 2800000

= 56000

2). variable manufacturing overhead = ( budgeted time for actual output * budgeted rate per hour) - (actual time * actual rate per hour)

= (56000 * rs 10) - ( 50400 * rs 13.5)

= (120400)

3). as variable manufacturing overhead is negative shows adverse situation and it is calculated with the help of actual output and hence are more reliable

24-  fixed budgeted manufacturing overhead rate = 4

fixed actual manufacturing overhead = 272000

1). fixed manufacturing variance = ( budgeted time for actual output * budgeted rate per hour) - (actual time * actual rate per hour)

= ( 56000 * 4) - 272000

= (48000)

2). fixed overhead is negative hence underllocated by 48000

3). the result of both fixed and variable overhead is not favourable means the estimate of cost gone wrong ans in actual more cost incurred which gives negative variance hence this driven that there is unefficiant estimate by the management and need to controll actual cost efficiently and effectively.